OpenStudy (anonymous):

In the expression (x+y)^8, if x=0.3 and y=0.7, what is the numerical value of the third term? 0.010, 0.020, 0.058, 0.060, or 0.068

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@phi @TheSmartOne

OpenStudy (anonymous):

@thomaster @confluxepic

OpenStudy (phi):

I assume using the binomial expansion ?

OpenStudy (anonymous):

@AravindG

OpenStudy (phi):

use \[\left(\begin{matrix}n \\ k\end{matrix}\right)a^{n-k} \ b^k\] where n is 8, "a" is 0.3 and "b" is 0.7 to find k for the third term start counting from 0 0 1 2 so k=2 is the third term. fill in the formula \[\left(\begin{matrix}8 \\ 2\end{matrix}\right)0.3^{8-2} \ 0.7^2 \\ 28 \cdot 0.3^6 \cdot 0.7^2 \\ 28 \cdot 0.000729 \cdot 0.49 \]